A random variable X is normally distributed with a mean of 121 and a variance of 121, and a random variable Y is normally distributed with a mean of 150 and a variance of 225. The random variables have a correlation coefficient equal to 0.5. Find the mean and variance of the random variable below. Av-218 (Type an integer or a decimal.) σ (Type an integer or a decimal.)
Answer #1

E(8)X + (-5)Y) = (8)E(X) + (-5)E(Y) = (8)(121) + (-5)(150) V((8)X + (-5)y) = V((8)X) + V((-5)y) (using independence) (8)-V(X)